Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson

Too cloudy for twilight sky flats. 0.6" seeing.

Taking test images with Br-gamma filter -- detector looks saturated. Restarting GEIRS server fixed the problem.

7:00 pm (2:00 UT) start telluric acquisition for MACS 0717 (Rex). Patchy ciurrus, 0.6" seeing.

7:30 pm (2:30 UT) start 2.5 hour K-band science observations for MACS 0717. Clearing, 0.5" seeing.

10:30 pm (5:30 UT) calibration and telluric observations for MACS 0717.

11:00 pm (6:00 UT) start 2.5 hour J-band science observations for MACS 0717. Clear, 0.6" seeing.

1:30 pm (8:30 UT) wind and humidity picking up. Seeing .8"--1".

2:00 pm (9:00 UT) start 70 minute re-observation of AEGIS 1418 (Rujopakarn) with new orientation. Clear, 0.8" seeing.

2:20 pm (9:20 UT) stopped tracking, power problem with the mount.

2:30 pm (9:30 UT) restarting acquisition for AEGIS 1418.

3:00 am (10:00 UT) same problem as before, stopped tracking, power problem with the mount.

3:03 am (10:03 UT) fixed again, resuming AEGIS 1418 observation (have to reacquire mask alignment).

3:11 am (10:11 UT) problem occurred again. Woke up Jeff (telescope manager on duty).

3:24 am (10:03 UT) trying a new fix, resuming AEGIS 1418 observation (have to reacquire mask alignment).

3:45 am (10:45 UT) seeing ~ 1".

4:20 am (11:20 UT) start observations of Abell 2218a (Rujopakarn). Clear, seeing 0.7".

6:00 am (13:00 UT) stopped wavefront sensing, running final calibration dark scripts.

Another MOS error while putting mask 6 (longslit) in storage. Will let German team debug it and run darks tomorrow afternoon.
Topic revision: r3 - 11 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback